

********************************************************************
UARS WEEKLY STATUS REPORT - 18 FEB 2005
********************************************************************
***********
SOLSTICE
***********
SOLSTICE has been under daylight only operations for four days this
week. Normal operations resumed Monday (doy 45). The instrument
continues to perform well, making the standard set of solar
observations daily.
***********
SUSIM
***********
SUSIM was turned off on Feb. 11 for the UARS yaw-around and resumed
operations on Feb. 14. Since then, the experiment successfully gathered
daily mid (1.1 nm) resolution, MgII, occultation, solar pointing offset,
and continuous monitoring scans for time periods with UARS telemetry or
tape recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004279 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004280 and 2004289 are currently waiting archival at GSFC DAAC.
HEPSA data between 2004290 and 2004299 are currently being processed.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004279. MEPS data between 2004280 and 2004299
are currently waiting archival at GSFC DAAC.
PEM continues to receive NMC data files from NOAA. NMC host machine
changed during mid-January which caused some NMC data transfer failures;
however, this situation was corrected and recovery completed with the
assistance of Roger Lin at NOAA. No problems have been encountered since
updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 05 February 2005 (Orbit 73361, GMT Day 036) through
11 February 2005 (Orbit 73466, GMT Day 042).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 038/2319 73406 POWER OFF (YAW)
PEM 040/2314 73436 ZEPS/NEPS POWER OFF
SUSIM 041/2201 73450 POWER OFF
SOLSTICE 041/2312 73451 FULLTIME TO DAY ONLY
The beta angle ranged between the angles of 21.5 degrees to 05.7
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 1.0 degrees on 12 February 2005 (DOY 043).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.64 to 0.66 degrees C. Additional
battery performance data for each battery on 04 February 2005 and 11
February 2005 are provided in Appendix D.
The clock error ranged from 10.4 to -10.1 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 02
HALOE = 01
PEM = 01
SUSIM = 01
SOLSTICE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 06
Four (4) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the batteries have performed nominally and the MPS continues
to utilize the VT-7 charge rate.
The FOT has completed their CMS software acceptance testing. The next
step in the process is for the official release of the software to be
installed on the backup CMS. This will be done by the software
development team next week. Once that is accomplished, the FOT will
run a few simulated days of operations to confirm the install. If all
goes well, then the new release will be put onto the operational CMS.
MEETINGS
The FOT called a kickoff meeting on 07 February 2005 to start planning
the activities needed to reset the flight software epoch year onboard.
This will be done to prevent a clock rollover, which is scheduled to
occur on 16 September 2005.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 12 February 2005 (DOY 043).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 11 February 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
73406 038/2319 05-036 HALOE POWER OFF (YAW)
73422 040/0100 05-037 OBC CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
73436 040/2314 05-038 PEM ZEPS/NEPS POWER OFF
73441 041/0738 05-039 OBC CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
73450 041/2201 05-040 SUSIM POWER OFF
73451 041/2312 05-041 SOLSTICE FULLTIME TO DAY ONLY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 11 February 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 11 February 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
73422 040/0102 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
73441 041/0739 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 11 February 2005
04 February 2005 - GMT Day 035 Beta = 28.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -1.48/-2.42 +00.0/+00.0 07.0 0.00 0.0
2 -5.6/ -67.2 +10.06/+9.06 +28.4/-15.2 26.7 1.07 20.9
3 22.4/ -72.8 +7.74/+7.09 +26.8/-14.4 26.7 1.05 20.1
11 February 2005 - GMT Day 042 Beta = 05.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.42/-3.68 +00.0/+00.0 07.0 0.00 0.0
2 67.2/ 11.2 +8.40/+7.74 +28.0/-12.8 27.0 1.06 18.8
3 -33.6/-112.0 +5.79/+5.15 +25.6/-12.0 26.9 1.04 17.8
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 11 February 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-005 FOT 73380 037/0655 30832 Generic Late Acquisition
05-005 FOT 73388 037/1919 30832 Generic Late Acquisition
05-005 FOT 73413 039/1117 30832 Generic Late Acquisition
05-008 FOT 73431 040/1517 38482 IP NOC Router Problem (:37)
05-005 FOT 73431 040/1517 30832 Generic Late Acquisition (:42)
05-009 FOT 73448 041/1836 N/A Mutual Interference
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-006 WSGT 73303 032/0336 38406 WS not sending UPDs for
multip
05-007 FOT 73344 034/2039 N/A SUSIM RESET
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-005 FOT See List Feb 05 30832 Generic Late Acquisition
05-002 MPS 73008 012/1251 N/A BATT 3 Diff Voltage High
05-008 FOT 73431 040/1517 38482 IP NOC Router Problem (:37)
05-009 FOT 73448 041/1836 N/A Mutual Interference